OpenStudy (anonymous):

How do you simplify: 4/7 = x/119

OpenStudy (abb0t):

You can cross multiply.

OpenStudy (anonymous):

\[\frac{ 4 }{ 7 }\] +

OpenStudy (abb0t):

To get \((4 \times 119) = [7 \times x]\) then solve for x.